### v1

# Barking Up The Syntactic Tree: Enhancing VLM Training with Syntactic Losses

Abstract
Vision-Language Models (VLMs) achieved strong performance on a variety of tasks (e.g., image-text retrieval, visual question answering). However, most VLMs rely on coarse-grained image-caption pairs for alignment, relying on data volume to resolve ambiguities and ground linguistic concepts in images. The richer semantic and syntactic structure within text is largely overlooked. To address this, we propose HIerarchically STructured Learning (HIST) that enhances VLM training without any additional supervision, by hierarchically decomposing captions into the constituent Subject, Noun Phrases, and Composite Phrases. Entailment between these constituent components allows us to formulate additional regularization constraints on the VLM attention maps. Specifically, we introduce two novel loss functions: (1) Subject Loss, which aligns image content with the subject of corresponding phrase, acting as an entailment of standard contrastive/matching losses at the Phrase level; (2) Addition Loss, to balance attention across multiple objects. HIST is general, and can be applied to any VLM for which attention between vision and language can be computed; we illustrate its efficacy on BLIP and ALBEF. HIST outperforms baseline VLMs, achieving up to +9.8% improvement in visual grounding, +6.3% in multi-object referring segmentation, +1.1% in image-text retrieval, and +0.2% in visual question answering, underscoring the value of structuring learning in VLMs.
1 Introduction
Figure 1: Motivation. Existing VLM models train from image-caption pairs, requiring large corpora of data to establish word-region associations. The proposed HIerarchically STructured (HIST) learning framework, and associated losses, through syntactic analysis, decompose captions into hierarchy of noun phrases establishing entailment between them and the image and among the phrases themselves. Specifically, it extracts nouns (Subject level) from noun phrases and aligns them to the image, along with the corresponding noun phrases themselves (Noun Phrase level). In addition, it combines two noun phrases together into longer composite phrases and regularizes attention to be a sum of the constituent noun sub-phrases (Composite Combined Phrase level).
Vision-Language Models (VLMs) [37, 27, 25, 26, 44, 2, 18, 23, 29, 33, 15], pretrained on vast amounts of image-text pairs, generalize effectively across a wide variety of downstream tasks, including image-text retrieval [31, 28], visual question answering [11], and visual grounding [17, 46]. However, VLMs primarily leverage a direct, coarse-grained alignment between image and caption pairs, often overlooking the deeper structures embedded in the textual data.
Historically, the natural language processing (NLP) has made significant progress in understanding textual structure through the development of various parsing methods. Techniques such as syntactic parsing [30, 32], which decomposes sentences into hierarchical syntax trees, have been instrumental in identifying linguistic patterns and relationships between words [32]. These methods reveal the internal structure of text, allowing models to better capture the semantics within a sentence. Parsing has been effectively utilized in numerous tasks such as machine translation [3, 10] and sentence classification [20, 41], where understanding the hierarchical relationships in text provides substantial gains in accuracy. Syntactic parsing has also been shown to be useful in image-text retrieval [40], with order embeddings, and visual grounding [42]; preceding recent advances in LLMs and VLMs. However, such efforts have largely been abandoned more recently. Their efficacy and the general question of how richer textual representations can benefit VLM training remains open.
To this end, we propose a novel HIerarchically STructured (HIST) learning framework that strengthens the connection between vision and language by embedding hierarchical text structures into the training process, without requiring annotation beyond image-text pairs. Our framework decomposes captions into noun phrases, each focusing on a distinct object within the image, and organizes these phrases into a three-level hierarchy: Subject Level, Noun Phrase Level, and Composite Combined Phrase Level, as depicted in Figure 1. The Combined Phrase level serves as the root, formed by joining two object-centric phrases using an “and” [^1] We do this, instead of using the raw structure of the caption, in order to get richer syntactic structure potentially derived from multiple captions.. The object-centric phrases themselves are represented at the (middle) Noun Phrase level. Finally, the parsed subject of each noun phrase form the Subject level and the leaf of the syntactic hierarchy.
Entailment between these constituent components of the sentence and the image, allows us to formulate additional regularization constraints for training of VLMs. Specifically, we leverage three losses. At the Phrase Level we ensure alignment (entailment) between noun phrases and the image by leveraging standard VLM loses typically applied for image-caption pairs: image-text contrastive (ITC), image-text matching (ITM), and masked language modeling (MLM). At the Subject Level we similarly use ITC and ITM loses, but based on a single word noun. The motivation comes from the fact that a potentially much longer noun phrase, at the phrase level, may obfuscate text-image alignment by focusing more on adjectives (that maybe common among many phrases) as compared to the noun that tends to be more semantically important. The final, novel Addition loss, is formulated on emergent localization properties of VLMs [16, 24, 25] and requires the sum of the attention maps for the noun sub-phrases to be as close as possible to the combined composite phrase attention. This loss encourages the model to attend to multiple objects concurrently, rather than disproportionately focusing on the most prominent one. Collectively, these losses regularize training of VLMs, improving their performance, without requiring additional supervision beyond traditional image-caption pairs.
Contributions. Our contributions are two-fold:
  • Based on syntactic parsing, we propose a HIerarchically STructured weakly-supervised framework (HIST) that improved image-text alignment in VLM training.
  • We design two novel loss functions: (1) Subject Loss to ensure image alignment with both object-centric phrase and the subject, and (2) Addition Loss to encourage equal attention across multiple objects in a phrase.
Empirical results demonstrate the efficacy of our framework, showing an improvement of up to 9.8% on visual grounding and +6.3% on multi-object referring segmentation. Somewhat surprisingly, this approach, which was largely designed to improve spatial image-text alignment in VLMs, also benefits them more broadly with improvement of up to +1.1% on image-text retrieval, and +0.2% on visual question answering (VQA), underscoring the broader advantages of sentence structure modeling in VLMs.
Figure 2: The overall structure of HIST. We decompose image captions into object-centric phrases and build a three-level hierarchy – Subject level, Phrase level, and Composite Combined Phrase level. Entailment between these constituent components of the sentence and the image, allows us to formulate additional regularization constraints for training of VLMs. Specifically, we leverage three losses. At the Phrase Level we ensure alignment (entailment) between noun phrases and the image by leveraging standard VLM loses typically applied for image-caption pairs: image-text contrastive (ITC), image-text matching (ITM), and masked language modeling (MLM) or language modeling (LM), depending on the model. At the Subject Level we similarly use ITC and ITM loses, but focus on matching the image and the subject noun of the phrase. This prevents adjectives in the phrase from obscuring the text-image alignment. The final Addition loss requires the sum of the attention maps for the noun phrases to be as close as possible to the combined composite phrase attention. This loss encourages the model to attend to multiple objects concurrently, rather than disproportionately focusing on the most prominent one.
2 Related Work
Vision-Language Models (VLMs). Vision-language pre-training (VLP) has established itself as a cornerstone in multimodal learning, driving advancements across tasks. Various architectures have been developed to optimize these tasks, including dual-encoders like CLIP [33] and ALIGN [15] for scalable contrastive learning, fusion-encoders like ALBEF [24] and LXMERT [38] for early cross-modal integration, and unified transformers such as BLIP [25] and BEiT [4], which consolidate multiple capabilities into a single model with both encoder and decoder architecture. Recently, VLMs have also begun integrating large language models (LLMs), as seen in BLIP2 [26] and LLava [29], enhancing multimodal interactions by combining powerful language understanding with visual components. Key training objectives in VLP have converged around image-text contrastive learning [33, 24, 25, 23, 47, 37], image-text matching [24, 25, 44, 23, 47, 37, 18, 43], and (masked) language modeling [44, 23, 47, 37, 6, 23, 2, 25], enabling refined and versatile multimodal understanding.
Despite these strides, most VLMs rely on pairwise image-text matching, often overlooking hierarchical structures in data, such as relationships among sub-phrases. Addressing this, our approach introduces hierarchical alignment, regularizing VLM learning with these nuanced relationships. This method enhances the model’s ability to localize and ground language-described entities in images; without the need for additional annotations or data.
Vision-Language Learning with Language Structure Supervised Losses. Several prior works have explored the benefits of leveraging language structure to enhance vision-language understanding. Early studies, such as [40], demonstrated the advantage of encoding textual hierarchical relationships (like generalization and entailment) in embedding spaces to improve tasks such as hypernym prediction, image-caption retrieval, and textual entailment. Another work [42] applied compositionality and exclusivity constraints, across noun phrases, to attention maps to improve visual grounding. However, these works were limited to smaller models and datasets, leaving open the question of to what extent these findings hold for the recent VLMs.
More recently, [8] introduces a semi-supervised approach that aligns image regions with corresponding noun phrases in captions, improving vision-language models (VLMs) on tasks such as visual question answering, visual commonsense reasoning, and visual grounding. Related, [5] examines how attributes, such as color, size, and texture, within image captions influence the learning and performance of vision-language models on open-vocabulary object detection and fine-grained text-region retrieval. SelfEQ [14] fine-tunes VLMs to produce consistent visual explanations across paraphrased captions, showing marked improvements in visual grounding tasks.
Conceptually related [22] and [48] apply constraints to GradCAM maps according to linguistic structure, with the aim of enhancing visual grounding. In [48], dependency parsing on image captions is used to identify distinct objects and prevents overlapping attention maps between them, thereby improving model performance on visual grounding. [22] designs specific losses for referring segmentation, enforcing that words within the same noun phrase have similar GradCAM maps while suppressing attention on incorrect object regions based on relationships with neighboring words (building on observations in [42]).
We are inspired by [14] and [22], however, propose a formulation that models complementary sentence structure in terms of hierarchy of phrase and sub-phrase relations. As such, we posit that our loses and objectives can be combined with consistency [14] and exclusion [22] loses designed in those works. Moreover, we are the first, to our knowledge, to show that such regularization losses not only improve visual localization, but general VLM understanding.
3 Method
Our approach aims to improve the visual-textual alignment of VLMs by fine-tuning them with a novel HIerarchicaly STructured (HIST) set of losses for the constituent text phrases. Specifically, we decompose image captions into object-centric phrases and build a three-level hierarchy – Subject level, Phrase level, and Composite Combined Phrase level – as shown in Figure 1; with the steps specified in Algorithm 1 and Section 3.2. The hierarchical structure allows us to design training objectives that achieve more granular alignment between visual and textual features.
In Section 3.1, we first introduce preliminaries and the base model architectures to which we apply proposed HIST. These base models are typically optimized with three widely used losses formulated atop of image-text pairs: Image-Text Contrastive (ITC), Image-Text Matching (ITM), and Masked Language Modeling (MLM). We use these losses in conjunction with our hierarchically structured losses to fine-tune the base models. In Section 3.2, we present our hierarchical alignment objectives: the Subject loss and the Addition loss. The Subject loss aims to align the main subject of each noun phrase with its corresponding visual representation, enforcing consistency between subjects and corresponding phrases. The Addition loss further guides the model to attend to multiple objects within the same image by aggregating Grad-CAM maps across different phrases, ensuring that the model’s attention covers all relevant objects. This hierarchical loss structure allows our model to align fine-grained visual and textual details while maintaining global coherence across the image-text pair.
Experimentally, we show that these additional losses, that require no added supervision beyond image-text pairs, lead to significant improvement on grounding. Further, we show that they moderately improve performance on a variety of other downstream VLM tasks (see Sec. 4.3).
Algorithm 1 Data Pair Creation: Noun Phrases Generation and Subject Constraints
Input: Dataset $\mathcal{D}=\{(\mathbf{I}_{k},\text{cap}_{ik})\mid k=1,2,\dots,M,\,i=1,2,\dots,N\}$, where each image $\mathbf{I}_{k}$ has $N$ captions, $\text{cap}_{ik}$, used to generate object-centric phrases.
Result: Set of data pairs $\mathcal{P}$.
1:$\mathcal{P}\leftarrow\emptyset$triangleright Initialize an empty set of pairs
2:for each $(\mathbf{I}_{k},\text{cap}_{k})$ in $\mathcal{D}$ do
3:     $\{(\mathbf{C}_{1}^{k},S(\mathbf{C}_{1}^{k})),\dots,(\mathbf{C}_{N}^{k},S(\mathbf{C}_{N}^{k}))\}\leftarrow$ LLM$(\text{cap}_{k})$triangleright Generate noun phrases and their subject
4:     for each phrase $\mathbf{C}_{i}^{k}$ in $\{\mathbf{C}_{1}^{k},\dots,\mathbf{C}_{N}^{k}\}$ do
5:         $\text{Candidates}\leftarrow\{\mathbf{C}_{j}^{k}\mid j\neq i\land S(\mathbf{C}_{j}^{k})\neq S(\mathbf{C}_{i}^{k})\}$
6:         if $\text{Candidates}\neq\emptyset$ then
7:              Randomly select $\mathbf{C}_{j}^{k}$ from Candidates
8:              $\mathcal{P}\leftarrow\mathcal{P}\cup\{(\mathbf{I}_{k},\mathbf{C}_{i}^{k},\mathbf{C}_{j}^{k})\}$
9:         end if
10:     end for
11:end for
12:return $\mathcal{P}$
3.1 Base Model and Objectives
Our formulation, HIST, is applicable to any VLM model that is capable of obtaining spatial image-text attention. We illustrate its efficacy on two popular VLM models: ALBEF [24] and BLIP [25]. Both of these models consist of an unimodal image encoder $\boldsymbol{\phi_{v}}$, a unimodal text encoder $\boldsymbol{\phi_{t}}$, and a multimodal encoder $\boldsymbol{\phi_{m}}$. The output of the image encoder and text encoder are represented as $\mathbf{V}=\{\mathbf{v}_{cls},\mathbf{v}_{1},\cdots,\mathbf{v}_{N_{v}}\}\in\mathbb{R}^{(N_{v}+1)\times d_{v}}$, where $N_{v}$ is the number of visual tokens and $d_{v}$ is the visual embedding dimension, and $\mathbf{T}=\{\mathbf{t}_{cls},\mathbf{t}_{1},\cdots,\mathbf{t}_{N_{t}}\}\in\mathbb{R}^{(N_{t}+1)\times d_{t}}$ where $N_{t}$ is the number of text tokens and $d_{t}$ is the text embedding dimension. Both models utilize the three commonly used objectives for learning visual and textual representations: image-text contrastive (ITC), image-text matching (ITM), masked language modeling (MLM). We provide a brief overview of these objectives here as we also leverage them during our training.
Image-Text Contrastive Learning (ITC). The ITC objective aims to align unimodal representations by maximizing similarity for matching image-text pairs and minimizing it for non-matching pairs. For an image $\mathbf{I}$ and text $\mathbf{T}$, we define the similarity as:
$s(\mathbf{I},\mathbf{T})=g_{v}(\mathbf{v}_{\text{cls}})^{\top}g_{t}(\mathbf{t}_{\text{cls}}),$
where $g_{v}$ and $g_{t}$ are linear projections of the [CLS] token from the image patches embedding and the text embedding.
The image-to-text and text-to-image similarities are calculated as:
$p^{i2t}_{m}(\mathbf{I})=\frac{\exp(s(\mathbf{I},\mathbf{T}_{m})/\tau)}{\sum_{m=1}^{M}\exp(s(\mathbf{I},\mathbf{T}_{m})/\tau)},$
$p^{t2i}_{m}(\mathbf{T})=\frac{\exp(s(\mathbf{T},\mathbf{I}_{m})/\tau)}{\sum_{m=1}^{M}\exp(s(\mathbf{T},\mathbf{I}_{m})/\tau)},$
where $\tau$ is a temperature parameter and M being the image or text queue following MoCo [13]. The ITC loss is defined as follows:
$\mathcal{L}_{\text{itc}}=\frac{1}{2}\mathbb{E}_{(\mathbf{I},\mathbf{T})\sim\mathcal{D}}\Big{[}\mathcal{H}\big{(}y^{i2t}(\mathbf{I}),p^{i2t}(\mathbf{I})\big{)}+$
$\hskip 137.31255pt+\mathcal{H}\big{(}y^{t2i}(\mathbf{T}),p^{t2i}(\mathbf{T})\big{)}\Big{]},$
where $y^{i2t}(I)$ and $y^{t2i}(T)$ are one-hot ground-truth vectors, $\mathcal{H}$ denotes the cross-entropy, and $\mathbb{E}$ is expectation over the training dataset $\mathcal{D}$ in image-text pairs.
Image-Text Matching (ITM). The ITM objective predicts whether an image-text pair $(\mathbf{I},\mathbf{T})$ is positive (matched) or negative (not matched). This is achieved by passing the fused [CLS] embedding from the multimodal encoder through a fully connected layer followed by a softmax to predict a binary probability $p^{\text{itm}}(\mathbf{I},\mathbf{T})$. The ITM loss $L_{\text{itm}}$ is defined as:
$\mathcal{L}_{\text{itm}}=\mathbb{E}_{(\mathbf{I},\mathbf{T})\sim\mathcal{D}}\,\mathcal{H}\left(y^{\text{itm}},p^{\text{itm}}(\mathbf{I},\mathbf{T})\right),$
where $y^{\text{itm}}$ is a two-dimensional one-hot vector representing the ground-truth label (matched or not).
Masked Language Modeling (MLM). Masked Language Modeling (MLM) integrates contextual information from both the text and the paired image to predict masked tokens within the text. Tokens in the input text are randomly masked out and replaced with a special [MASK] token. Let $\hat{\mathbf{T}}$ denote the masked text and $p^{\text{msk}}(\mathbf{I},\hat{\mathbf{T}})$ denote the model’s predicted probability distribution for the masked token. The MLM objective $\mathcal{L}_{\text{mlm}}$ minimizes the cross-entropy loss:
$\mathcal{L}_{\text{mlm}}=\mathbb{E}_{(\mathbf{I},\hat{\mathbf{T}})\sim\mathcal{D}}\,\mathcal{H}\left(y^{\text{msk}},p^{\text{msk}}(\mathbf{I},\hat{\mathbf{T}})\right),$
where $y^{\text{msk}}$ is a one-hot distribution representing the ground-truth token for the masked position.
Certain VLMs though, like BLIP [25], use a language modeling loss focused on predicting the next word token rather than randomly masking words within sentences and predicting the masked tokens. Essentially, this approach uses a cross-entropy loss for next-word prediction.
3.2 HIerarchically STructured (HIST) Losses
To enhance cross-modal alignment at different levels of granularity, we introduce two additional sets of hierarchically structured losses which we describe in this section; mainly the Subject and the Addition loss.
Phrase parsing. As illustrated in Algorithm 1, we follow SelfEQ [14] and perform phrase chunking using Vicuna [7]. Specifically, in-context examples are provided to Vicuna [7] to first generate noun phrases from caption, and then each noun phrase is re-inputted to Vicuna to detect the subject. To construct composite phrase dataset $\mathcal{P}$, for each noun phrase $\mathbf{C}_{i}^{k}$, we identify potential pairing candidates $\{\mathbf{C}_{j}^{k}\}$ from other noun phrases from the same image but not referring to the same subject (these do not necessarily come from the same caption). We then randomly select a pairing candidate and form a data pair $(\mathbf{I}_{k},\mathbf{C}_{i}^{k},\mathbf{C}_{j}^{k})$, and add it to $\mathcal{P}$. This process repeats for all phrases in each caption, ultimately generating $\mathcal{P}$, which contains the selected image-composite caption phrase pairs for further use.
Phrase Loss. With the noun phrases we obtained above, we apply the ITC, ITM and MLM on the phrases directly: $\mathcal{L}_{phrase}=$
$\mathcal{L}_{ITC}(\{(\mathbf{I},{\mathbf{C}_{i}})\})+\mathcal{L}_{ITM}(\{(\mathbf{I},{\mathbf{C}_{i}})\})+\mathcal{L}_{MLM}(\{(\mathbf{C}_{i})\}).$
Subject Loss. The motivation of the Subject loss comes from the fact that a potentially much longer noun phrase, at the phrase level, may obfuscate text-image alignment by focusing more on adjectives (that maybe common among many phrases) as compared to the noun that tends to be more semantically important. With such an observation, we add a Subject Loss during training, which enforces consistency, aligning the subject of the phrase with the corresponding visual representation. Specifically, for a phrase $\mathbf{C}_{i}$ with main subject $\mathbf{S}_{i}$, the loss is:
$\mathcal{L}_{subject}=\mathcal{L}_{ITC}(\{(\mathbf{I},{\mathbf{S}_{i}})\})+\mathcal{L}_{ITM}(\{(\mathbf{I},{\mathbf{S}_{i}})\}).$
where $\mathcal{L}_{ITC}$ and $\mathcal{L}_{ITM}$ are defined above.
Localization in VLMs. We first introduce how we obtain localization map for phrases and then discuss the novel Addtion loss we proposed.
The text prompt and the image first go through the modality-specific encoders $\boldsymbol{\phi_{t}}$ and $\boldsymbol{\phi_{v}}$ respectively, then the respective embeddings are input into the cross-attention fusion module. For each cross-attention layer, the text encodings serve as query vectors and the image patch encodings are as key and value vectors. From a particular cross-attention layer (and specific head), we extract the corresponding text-to-image cross-attention map as shown below. Note that different layers and heads lead to drastic performance differences, and we follow [14, 16] for the best layer-head combination; resulting in
$\mathbf{A}_{\mathbf{C}_{i}}=\texttt{Softmax}\left(\frac{\boldsymbol{\phi_{t}}(\mathbf{T})\ \mathbf{W}_{Q}(\boldsymbol{\phi_{v}}(\mathbf{I})\mathbf{W}_{K})^{\top}}{\sqrt{D_{t}}}\right),$
where $\mathbf{A}_{\mathbf{C}_{i}}$ has shape $(K+1)\times P\times P$ with $P$ being the number of image patches and $K$ being the number of text tokens. To get the phrase specific attention map, we take the mean over all K tokens and get final map of size $P\times P$. This map can serve as proxy for segmentation [16] and can further be enhanced by GradCAM [34], to highlight discriminative regions, as shown in [39, 16]. Applying this technique results in the final localization map:
$\mathbf{G}_{\mathbf{C}_{i}}=\texttt{ReLU}\left(\mathbf{A}_{\mathbf{C}_{i}}\cdot\frac{\partial\mathcal{L}_{\text{ITM},i}}{\partial\mathbf{A}_{\mathbf{C}_{i}}}\right),$
for a given phrase $\mathbf{C}_{i}$ and a particularly chosen attention layer and head. This final localization map is obtained by taking the element-wise product of the cross-attention map with a gradient of the loss $\mathcal{L}_{\text{ITM},i}$ with respect to it.
Addition Loss. We are now ready to define our novel Addition loss at the combined phrase level. For a pair of phrases $\mathbf{C}_{i}$, $\mathbf{C}_{j}$, extracted from captions of the same image, and the combined composite phrase $\mathbf{C}_{ij}$ we extract the cross-attention maps and then the corresponding localization maps as follows:
$\mathbf{G}_{\mathbf{C}_{i}}=\texttt{ReLU}\left(\mathbf{A}_{\mathbf{C}_{i}}\cdot\frac{\partial\mathcal{L}_{\text{ITM},i}}{\partial\mathbf{A}_{\mathbf{C}_{i}}}\right),$
$\mathbf{G}_{\mathbf{C}_{j}}=\texttt{ReLU}\left(\mathbf{A}_{\mathbf{C}_{j}}\cdot\frac{\partial\mathcal{L}_{\text{ITM},j}}{\partial\mathbf{A}_{\mathbf{C}_{j}}}\right),$
$\mathbf{G}_{\mathbf{C}_{ij}}=\texttt{ReLU}\left(\mathbf{A}_{\mathbf{C}_{ij}}\cdot\frac{\partial\mathcal{L}_{\text{ITM},ij}}{\partial\mathbf{A}_{\mathbf{C}_{ij}}}\right).$
We observed that vision-language models often focus disproportionately on the most prominent objects within an image, which can lead to poor alignment. Therefore, we introduce the Addition loss to improve the model’s capacity to attend to multiple objects simultaneously. It is defined as:
$\mathcal{L}_{\text{add}}=\sum_{P}\sum_{P}\left|\mathbf{G}_{\mathbf{C}_{i}}+\mathbf{G}_{\mathbf{C}_{j}}-\mathbf{G}_{\mathbf{C}_{ij}}\right|.$
The loss enforces that the attention map for the combined phrase approximates the sum of individual attention maps, ensuring that the model considers all relevant objects.
3.3 Total Hierarchical Alignment Loss
The total loss combines all the losses mentioned above:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{phrase}}+\mathcal{L}_{\text{subject}}+\mathcal{L}_{\text{add}}.$
This hierarchical loss formulation allows our model to align fine-grained visual and textual semantics effectively, while maintaining global coherence across the image-text pair. Each component contributes to a specific level of alignment, from individual phrases to subjects and combined phrases, ultimately refining the model’s ability to handle complex, multi-object scenes and improving its localization ability.
4 Experiment
Implementation details. We apply HIST on ALBEF [24] and BLIP [25] and train the model with Visual Genome [21] and MSCOCO [28]. ALBEF combines a ViT-B model for encoding images and a BERTbase model for encoding text. Following SelfEQ [14], we train ALBEF with image size of $256\times 256$ with the compostie objective function $\mathcal{L}$ = $\alpha$ $\mathcal{L}_{\text{vl}}$ + (1 - $\alpha$) ($\mathcal{L}_{\text{vl}}$ + $\mathcal{L}_{\text{HIST}}$), where $\alpha$ is initially set to 0 and increments to 1, remaining constant after the second epoch. We train our model with with 4 NVIDIA A6000 with learning rate of 2e-5. For the addtion loss, we employ the head average gradcam map from the 3rd cross-attention layer. For BLIP, we train BLIPlarge, which adopts VIT-L/16 as the image encoder, BERT as the text encoder, and insert an extra cross-attention layer for each transformer block of BERT. Following BLIP [25], we train BLIP with image size of $224\times 224$, on 4 NVIDIA A6000 with learning rate 1e-5. For the addtion loss, we leverage the head average gradcam map from the 9th cross-attention layer. We will release all models, data and code upon acceptance of the paper.
4.1 Visual Grounding
Method Training Flickr30K Referit
InfoGround [12] VG-boxes 76.7 -
VMRM [9] VG-boxes 81.1 -
AMC [45] VG-boxes 86.6 73.2
ALBEF [24] Zero-shot 79.4 59.7
BLIP [25] Zero-shot 80.1 51.6
[36] VG 75.6 66.0
g++ [35] VG 80.0 70.3
ALBEF + SelfEQ [14] VG 81.9 67.4
(Ours) ALBEF + HIST VG 83.6 69.5
(Ours) BLIP + HIST VG 81.5 58.1
[36] COCO 75.4 61.0
g++ [35] COCO 78.1 61.5
ALBEF + SelfEQ [14] COCO 84.1 62.8
(Ours) ALBEF + HIST COCO 85.3 63.4
(Ours) BLIP + HIST COCO 84.7 57.6
Table 1: Visual Grounding on Flickr30K [31] and Referit [17]. VG stands for Visual Genome. Our HIST atop of ALBEF and BLIP outperform baselines and SelfEQ [14] variants.
Method Training RefCOCO+ Test A RefCOCO+ Test B
InfoGround [12] VG-boxes 39.8 44.1
VMRM [9] VG-boxes 58.9 50.3
AMC [45] VG-boxes 80.3 64.6
ALBEF [24] Zero-shot 69.4 53.8
BLIP [25] Zero-shot 70.2 52.4
ALBEF + SelfEQ [14] VG 74.1 55.1
ALBEF + SelfEQ [14] VG 75.1 55.5
(Ours) ALBEF + HIST VG 74.7 56.7
(Ours) BLIP + HIST VG 75.3 56.4
Table 2: Visual Grounding on RefCOCO+ [46]. VG stands for Visual Genome. represent our rerun result.
We evaluate ALBEF [24] and BLIP [25] on the test splits of Flickr30k [31] and ReferIt [17], which contain 1k and 10k images, respectively, following [14, 1]. Additionally, following SelfEQ [14], we assess our method on RefCOCO+ [46] Test A and Test B, a challenging benchmark for box-supervised methods. We use pointing game accuracy [49] for evaluation. It measures the percentage of test pairs where the highest point in the predicted heatmap for each pair falls within the ground-truth bounding box.
As observed in Table 1, when applying HIST to the base model ALBEF, we achieve an absolute performance gain of 4.2% on Flickr30K and 9.8% on ReferIt when trained with Visual Genome [21], and gains of 5.9% on Flickr30K and 3.7% on ReferIt when trained with MSCOCO [28]. Similarly, applying HIST to the base model BLIP yields a performance improvement of 1.4% on Flickr30K and 6.5% on ReferIt with Visual Genome [21], and improvement of 4.6% on Flickr30K and 6.0% on ReferIt with MSCOCO [28].
Notably, we also outperform the state-of-the-art in Visual Grounding. Compared to the recent approach SelfEQ [14], which uses the same backbone and training data, we achieve an improvement of 1.7% on Flickr30k and 2.1% on ReferIt when trained on Visual Genome, and 1.2% on Flickr30k and 0.6% on ReferIt when trained on MSCOCO. Our method also performs comparably to methods using bounding box supervision on Flickr30k. Although our performance for visual grounding on Referit is marginally lower than g++ [35], it is worth noting that their approach benefits from additional supervision through the use of heatmaps from CLIP as pseudo-labels; Our method can potentially benefit from similar supervision.
Following SelfEQ [14], we also evaluate our method (trained with VG) on RefCOCO+ [46], a more rigorous benchmark, typically used to assess box-supervised techniques, with 750 images in each of the Test A and Test B sets. As shown in Table 2, we outperform the base model ALBEF by 5.3% and 2.9% on Test A and Test B respectively. Additionally, we achieve a 0.6% improvement on RefCOCO+ Test A and a 1.6% improvement on RefCOCO+ Test B compared to SelfEQ. Applying HIST with BLIP, we obtain a substantial improvement of 5.3% / 4.4% on RefCOCO+ Test A / Test B over zero-shot evaluation.
We could not directly compare with CAFT [48] as we were unable to replicate the baseline performance reported in their work [^2] ALBEF’s performance on Flickr30K is reported as 84.49 in [48], while SelfEQ [14] and our experiments find it to be 79.4.. However, our improvements over the ALBEF baseline are relatively larger compared to CAFT. Notably, our losses and objectives are different and complementary to both SelfEQ [14] and CAFT [48], allowing our method to be integrated with their losses for potential gains.
Figure 3: Qualitative Result for ALBEF, SelfEQ and HIST + ALBEF on Visual Grounding and Referring Segmentation. Images are from RefCOCO+. The enlarged red stars represent the top 4 locations with the highest predicted attention value from respective methods. To obtain segmentation mask, we input the four points into SAM [19] as point prompts. We note that the HIST accurately detect all objects but the mask can over-segment or under-segment.
Method Training CIoU
ALBEF [24] Zero-shot 29.4
ALBEF + SelfEQ [14] COCO 30.3
(Ours) ALBEF + HIST COCO 35.6
Table 3: Complex Referring Segmentation on RefCOCO+ [46].
4.2 Complex Referring Segmentation
For Referring Segmentation, we focus on the RefCOCO+ [46] dataset to evaluate HIST’s performance in multi-object scenarios. We developed a custom multi-object referring segmentation test set from RefCOCO+ [46]. For each image in the RefCOCO+ test set that contains two or more text phrases, we randomly select two phrases and combine them using an “and ”. This yields 1,404 unique image-text pairs designed to assess the ability to handle complex multi-object references.
To isolate the model’s localization ability from segmentation mask completeness, we utilize SAM [19] as a post-processing tool. For each method, we first get the location of the top $4$ patches in the $P\times P$ predicted heatmap, then we scale it to the corresponding location in the original image dimension. These points serve as the input to SAM to obtain corresponding object masks. We report CIoU following standard referring segmentation evaluation metrics [51, 50].
As shown in Table 3, we outperform both ALBEF and SelfEQ [14] by 6.2% and 5.3%, respectively, on the complex scene referring segmentation. Additionally, Figure 3 shows qualitative comparisons among ALBEF, ALBEF + SelfEQ, and ALBEF + HIST. We observe that ALBEF + HIST significantly outperforms ALBEF and ALBEF + SelfEQ in indentifying complex queries involving multiple objects within intricate scenes. For instance, it accurately identifies the “Blurry pie” in a blurry image, the occluded “Not full zebra” and relative object positions such as “The further meter,” the “Nearest meter,” and the “Lady in the back.” This enhanced ability to distinguish multiple objects in challenging contexts underscores HIST’s robustness and effectiveness in visual grounding tasks. (More qualitative result provided in the supplementary.)
4.3 Downstream Tasks
While our losses were designed to improve localization, we also test the impact of they have on other downstream VLM tasks, including image-text retrieval and VQA.
Image-Text Retrieval. For Image-Text Retrieval, we use the MSCOCO [28] and Flickr30k [31] datasets. We evaluate ALBEF + HIST (trained with VG) for both image-to-text retrieval (TR) and text-to-image retrieval (IR). Following ALBEF [24], we finetune the pre-trained model using ITC and ITM losses and the training samples from the respective datasets. We report average recall@1 for this task. From Table 4, we observe that compared to ALBEF, ALBEF + HIST gives 1.1% improvement on image-to-text retrieval recall@1 and 0.7% on text-to-image retrieval recall@1 when finetune on COCO, and 0.2% improvement on image-to-text retrieval recall@1 and 0.4% on text-to-image retrieval recall@1 when finetune on Flickr30k.
Retrieval-FT COCO Retrieval-FT Flickr30k
Method TR@1 IR@1 TR@1 IR@1
ALBEF [24] 77.6 60.7 95.9 85.6
(Ours) ALBEF + HIST 78.7 61.4 96.1 86.0
Table 4: Image-Text Retrieval on COCO [28] and Flickr30K [31]. We train ALBEF with HIST on Visual Genome.
Visual Question Answering. For VQA, we use the VQA [11] dataset. Following ALBEF, we fine-tune a 6-layer transformer decoder to generate answers, initializing it with pre-trained weights from the multimodal encoder of HIST + ALBEF. As shown in Table 5, We observe +0.2% improvement on VQA-dev and VQA-std.
Method VQA-dev VQA-std
ALBEF [24] 75.8 76.0
(Ours) ALBEF + HIST 76.0 76.2
Table 5: Visual Question Answering on VQA [11]. We train ALBEF with HIST on Visual Genome.
4.4 Ablation Study
HIST Losses. We ablate different loss components of HIST + ALBEF and report the results in Tab. 6. Each component contributes positively to the final performance, which highlights the significance of our design choices.
Data Objective Flickr30k
- $\mathcal{L}_{\text{vl}}$ 79.4
$T$ $\mathcal{L}_{\text{phrase}}$ 81.6
$T_{HIST}$ $\mathcal{L}_{\text{phrase}}$ + $\mathcal{L}_{\text{add}}$ 85.0
$T_{HIST}$ $\mathcal{L}_{\text{phrase}}$ + $\mathcal{L}_{\text{subj}}$ + $\mathcal{L}_{\text{add}}$ 85.3
Table 6: Ablation on HIST Losses. Models are trained with COCO. T represents training with only noun phrases of COCO. $T_{HIST}$ represents training with our HIST structure.
BLIP Layer and Head effect. As shown in Table 7, our experiments demonstrate that when training BLIP with HIST, using the head average attention map from the 9th cross-attention layer to calculate the Addition Loss enhances grounding performance in both layer 9 head 10, and layer 8 head 10 of BLIP. Consequently, we report BLIP’s visual grounding performance by training with the Addition loss using the head average Grad-CAM map from the 9th cross-attention layer and evaluate using the mean of two attention maps: (1) layer 9 head 10 and (2) layer 8 head 10.
Eval Train L8 L9 L8H10 L9H10 Mean Mul
L8H10 51.1 56.9 42.9 56.5 48.9 56.5
L9H10 48.3 51.8 53.9 32.4 17.1 32.4
Mean 51.2 57.6 53.9 56.7 18.7 56.8
Mul 52.4 56.2 53.7 49.6 21.4 49.6
Table 7: Ablation of HIST + BLIP on layer and head. We train HIST+BLIP on COCO and test on Referit with image size of 384. Mean represent the average of attention maps from L8H10 and L9H10 of BLIP. Mul represent the element-wise multiplication of attention maps from L8H10 and L9H10 of BLIP.
5 Conclusion
We present a novel HIerarchically STructured (HIST) learning framework that enhances vision-language alignment in Vision-Language Models (VLMs) by embedding hierarchical syntactic structures into training. HIST organizes captions into structured noun phrases across multiple syntactic levels, improving alignment at both phrase and object levels without additional supervision. Our framework introduces two key losses: Subject Loss, ensuring alignment with object-centric phrases and subjects, and Addition Loss, promoting balanced attention across multiple objects in complex scenes. HIST significantly enhances localization in challenging multi-object scenes and improves performance in tasks like image-text retrieval and visual question answering, underscoring the value of hierarchical structuring for spatial and semantic alignment in VLMs.